OpenStudy (anonymous):

find the exact value of cos(2tan^(-1)(20/21))

OpenStudy (anonymous):

\[\cos(2x)=2\cos^2(x)-1\] if i recall correctly, so you only need \(\cos(\tan^{-1}(\frac{20}{21}))\)

OpenStudy (anonymous):



OpenStudy (anonymous):

which means you only need h, which you find by pythagoras as \[h=\sqrt{20^2+21^2}=\sqrt{841}\] and so \(\cos(\tan^{-1}(x))=\frac{21}{\sqrt{821}}\)

OpenStudy (anonymous):

then use the formula above

OpenStudy (anonymous):

2cos^2(21/29)-1=cos2x how do you simplify that

OpenStudy (anonymous):

oh no hold on

OpenStudy (anonymous):

it is not cosine of 21/29 cosine IS 21/29

OpenStudy (anonymous):

it is \[\cos(\tan^{-1}(\frac{20}{21}))=\frac{21}{29}\]

OpenStudy (anonymous):

so \[\cos(2\tan^{-1}(\frac{20}{21})=2\times (\frac{21}{29})^2-1\]

OpenStudy (anonymous):

the 21/29 is the output, not the input

OpenStudy (anonymous):

so the final answer is 41/841?

OpenStudy (anonymous):

it worked! thanks so much. you're a life saver

OpenStudy (anonymous):

do you know how to do this problem? find all the solutions in the interval [0,2pi), rounded to 5 decimal places (enter all answers as a comma separated list) 5cos(2pi)=4

OpenStudy (anonymous):

makes no sense since you have no variable there

OpenStudy (anonymous):

\[5\cos(\pi x)=4\] \[\cos(\pi x)=\frac{4}{5}\] \[\pi x=\cos^{-1}(\frac{4}{5})\] might work if that is where the variable is

OpenStudy (anonymous):

correction..

OpenStudy (anonymous):

5cos(2THETA)=4